OpenStudy (anonymous):

The value 3pi/16 is a solution for the equation 2cos^2(4x)-1=0. A. True B. False

OpenStudy (anonymous):

true... 2cos^2(4*3pi/16) - 1 = 2cos^2(3pi/4) - 1 = 2(-sqrt2/2)^2 - 1 = 2(1/2) - 1 = 1 - 1 = 0

OpenStudy (anonymous):

maybe i'll just put down the solution as "calculator" the next time... :)